When x number of units are sold, the price of each unit (in dollars) is given by p= x/2 + 75

Find the unit price when the following quantities are sold: 2, 7, 9, 11.

Answer by checkley71(8403)

to solve just substitute 2,7,9 & 11 for x thus:
p=2/2+75
p=1+75
p=76 solution
--------------------------
p=7/2+75
p=3.5+75
p=78.5 solution
--------------------------
p=9/2+75
p=4.5+75
p=79.5 solution
-------------------------
p=11/2+75
p=5.5+75
p=80.5 solution
-----------------------